# On the Feasibility of Hijacking MLLMs’ Decision Chain via One Perturbation

### v1

@@ Abstract
Conventional adversarial attacks focus on manipulating a single decision of neural networks. However, real-world models often operate in a sequence of decisions, where an isolated mistake can be easily corrected, but cascading errors can lead to severe risks.
This paper reveals a novel threat: a single perturbation can hijack the whole decision chain. We demonstrate the feasibility of manipulating a model’s outputs toward multiple, predefined outcomes, such as simultaneously misclassifying “non-motorized lane” signs as “motorized lane” and “pedestrian” as “plastic bag”.
To expose this threat, we introduce Semantic-Aware Universal Perturbations (SAUPs), which induce varied outcomes based on the semantics of the inputs. We overcome optimization challenges by developing an effective algorithm, which searches for perturbations in normalized space with a semantic separation strategy. To evaluate the practical threat of SAUPs, we present RIST, a new real-world image dataset with fine-grained semantic annotations. Extensive experiments on three multimodal large language models demonstrate their vulnerability, achieving a 70% attack success rate when controlling five distinct targets using just an adversarial frame.
@@ 1 Introduction
Visual multimodal large language models (MLLMs) like Llava [22], Qwen-VL [2], GPT-4V [1], and Gemini [8] are susceptible to adversarial perturbations [28, 15], where image hijacks [4] enable precise control over models’ outputs, posing a significant threat to downstream content.
In real-world applications like autonomous driving [38] and robotic tasks [18], continuous image frames are constantly captured and fed into the MLLM for a sequence of decisions (i.e., decision chain). In such applications, an isolated mistake can be easily corrected, but cascading errors can lead to severe risks. However, current research on universal adversarial perturbations (UAPs) [26, 27] overlooks the need to mislead the model’s decision chain. Existing studies on MLLM even struggle to achieve universal image hijacking; they either construct perturbations per image [4, 24] or manipulate both images and prompts [23, 12].
Figure 1: Illustration of hijacking of the MLLM’s decision chain. The adversary pastes the perturbation onto the camera. Once the vehicle starts and perceives images, the MLLM accurately outputs the attacker’s predefined target content based on the image semantics, cumulatively guiding the vehicle to a predefined destination.
This paper aims to reveal a novel threat: the adversary can hijack the model’s decision chain via one perturbation. As illustrated in Figure 1, with this capability, the adversary only needs to generate and paste a noise frame before the car starts. When perceiving perturbed images, the system will accurately output the attacker’s predefined target content based on the image semantics, which cumulatively guides the vehicle to a predefined destination.
We refer to such perturbations as Semantic-Aware Universal Perturbations (SAUPs) due to their unique semantic-aware ability: they can “actively” sense the input image’s semantics and map it to a predefined label. SAUPs also maintain the universality property, enabling the perturbation to generalize across unseen inputs.
Searching for SAUPs presents new optimization challenges: supporting multiple distinct targets constrains the feasible solution space, and interference among targets leads to misalignment with the semantics. To overcome optimization challenges, we propose the Semantic-ORiented aTtack (SORT), which comprises two key components: normalized space optimization and semantic separation optimization. (1) Instead of using the raw pixel space like previous work, we leverage the space after image normalization as a stable foundation for optimization. (2) To address the obfuscation between predefined targets, we combine cross-entropy loss and margin loss.
For a comprehensive understanding of the semantic-aware ability of SAUPs, we not only use the existing image classification dataset (i.e., ImageNet [9]), but also annotate a new dataset called Real Image Sequence Trajectories (RIST). RIST contains more than 1,000 real-world images and comprises 28 distinct image trajectories, sourced from established autonomous driving [38] and robotic manipulation [25] datasets. RIST captures diverse vehicular navigation and indoor robotic tasks across varied environmental conditions, covering rich content.
@@
Extensive experiments are conducted on three MLLMs, Llava1.5, Qwen2.5-VL, and InternVL3, leveraging the ImageNet and RIST datasets to investigate performance across various semantic granularities. The results show that SAUPs easily hijack both models under various settings. For example, one perturbation is sufficient to mislead Qwen, achieving success rates of 93%, 77%, 61%, and 66% under 2, 3, 4, and 5 targets, respectively.
Our further analysis reveals that the effectiveness of SAUPs primarily lies in aligning image semantic directions with distinct target labels. We demonstrate that incorporating optimization in the normalized space is crucial for generating effective SAUPs, while employing semantic separation optimization can further enhance performance.  [^1] The code will be released upon the acceptance of this paper.
We summarize the main contributions as follows:
  • We reveal a novel threat that a perturbation can precisely hijack the MLLM’s decision chain, and analyze the perturbed image features in the latent space.
@@ 2.1 Adversarial Perturbations
Adversarial perturbations have been demonstrated [32, 14] that subtle perturbations added to input data could significantly alter model predictions. A pivotal advancement is the discovery of UAP [26, 36], which are input-agnostic perturbations effective across diverse samples.
However, current adversarial perturbation methods primarily focus on manipulating a single decision of neural networks, either causing misclassification (i.e., non-targeted attacks [6]) or steering toward a predefined target label (i.e., targeted attacks [27]). Real-world AI systems, such as autonomous driving or the embodied AI [18], often rely on sequential decision-making processes. This reduces the practical threat of existing attacks, as they cannot simultaneously mislead an entire decision chain.
@@ 2.2 Vulnerability of MLLMs
The rapid development of MLLMs [22, 2] has raised new security concerns. Recent studies [28, 30, 40, 39] reveal that MLLMs inherit the vulnerability to adversarial perturbations. According to the degree of control over the attack, existing methods can be categorized into two types: semantic misleading and image hijacking.
Semantic misleading [41, 21, 37] only requires the model to misunderstand the meaning of the image. They typically deceive the vision encoder of MLLMs, causing it to provide incorrect image representations to the downstream language model, and consequently leading to incorrect outcomes in subsequent visual tasks.
In contrast, image hijacking [24, 23, 4] aims for precise control over the model’s outputs. They often employ end-to-end optimization. Due to their objective to fully dictate the output, such attacks pose a greater threat. For example, attackers can induce the model to generate specific bash commands, or bypass rule checks in downstream tasks [4].
2.3 Distinction from Prior Work
Some previous studies have explored similar properties of adversarial perturbations, but they significantly differ from this paper. C-PGC [12] proposes a UAP against vision-language models, which requires manipulating both image and text modalities, and consequently, it cannot achieve image hijacks or attack the decision chain. MultiAttack [13] shows that a single perturbation can map various inputs to targets. However, this perturbation is not universal, being effective only within the training set and failing to expose the semantic-aware capability.
3 Methodology
@@
3.1 Problem Statement
Consider the MLLM as a function $f:\mathcal{V}\times\mathcal{T}\rightarrow\mathcal{Y}$, which maps an image $\mathbf{x}\in\mathcal{V}$ and a prompt $\mathbf{p}\in\mathcal{T}$ to a logit output $\mathbf{y}\in\mathcal{Y}$. The output $\mathbf{y}$ typically consists of a sequence of token logits produced autoregressively.
Our goal is to obtain a single adversarial perturbation $\delta\in\mathcal{V}$, that can be applied to any image $\mathbf{x}^{\left(c\right)}$ from the semantic clustering $c\in\left\{1,2,\dots,C\right\}$, causing the model to generate a predefined, class-specific target $\mathbf{t}^{\left(c\right)}=t^{(c)}_{1:N}$.
We refer to this type of perturbation as semantic-aware universal perturbations, characterized by its ability to induce different targets depending on the input semantics.
The workflow for SAUPs is illustrated in Figure 2. Formally, given a dataset $\mathcal{D}$ where each item consists of triples $\left\{\mathbf{p},\mathbf{x}^{(c)},\mathbf{t}^{(c)}\right\}$, the SAUP is obtained by solving the following optimization problem:
$\arg\min_{\delta}\;\mathbb{E}_{\mathcal{D}}\left\{\mathcal{L}\left(f\left(\mathcal{A}(\mathbf{x}^{(c)},\delta),\mathbf{p}\right),\mathbf{t}^{(c)}\right)\right\}$ (1)
where $\mathcal{L}$ is a loss function (details in Section 3.3), $\mathcal{A}\left(\cdot,\cdot\right)$ represents the perturbation append operation.
$\mathcal{A}(\mathbf{x},\delta)=\left[\mathbf{x}\cdot\left(\mathbf{1}-\mathbf{r}\right)+\delta\cdot\mathbf{r}\right]_{\mathbf{0}}^{\mathbf{1}}$ (2)
where $\left[\cdot\right]_{a}^{b}=\min\left(\max\left(\cdot,a\right),b\right)$, and $\mathbf{r}$ masks the allowable region where the perturbation can be applied.
Semantic Granularity and Real-World Dataset. Since SAUPs leverage the image semantics for mapping, it is necessary to clarify the semantic granularity. Existing research on adversarial perturbations [31, 4] and image classification [16, 9] focuses on the salient feature of the image (e.g., distinguishing between cats and dogs), which corresponds to coarse-grained semantics. In contrast, images can also be categorized according to fine-grained semantics (e.g., the specific dog apart from other dogs).
To bridge this gap and provide a more comprehensive assessment of the flexible threats posed by SAUPs, we propose Real Image Sequence Trajectories (RIST), an image dataset annotated with fine-grained semantics. RIST comprises over 1,000 images and 28 trajectories, collected from real-world videos and encompassing two representative scenarios: autonomous driving and robot tasking. As shown in Figure 3, RIST defines each class as a sequence of semantically coherent images, and each trajectory consists of multiple such classes extracted from a single video. Further details can be found in the Appendix.
By utilizing fine-grained semantic granularity, SAUPs enable attackers to flexibly orchestrate events. For instance, in the context of vision-based robotics, an image containing a knife (target action: hold) and an image featuring both a knife and a human (target action: throw) can be considered distinct semantic classes.
Figure 3: Illustration of RIST, which consists of real-world image trajectories across two scenarios: AutoDriving and RoboTasking. RIST clusters semantically similar frames as the same class, with each class containing 10 images.
3.2 Normalized Space Optimization
To facilitate the addition of perturbations to images, existing studies [4, 23, 35, 34] conduct their optimization directly in the pixel space ([0, 1]). While this approach is sufficient for existing attack objectives, it is inadequate for the stringent restrictions required by SAUP. Consequently, we are motivated to shift the optimization process to a more effective representation space, where the search for viable perturbations becomes more tractable.
We refer to the space after image normalization as an effective domain for searching SAUPs for the following reasons: 1) The loss landscape in the normalized space is smoother [19], resulting in stable gradients and fewer local optima. This benefits gradient-based optimization methods, making it easier to generate feasible perturbations. 2) Representations in the normalized space tend to have similar scales, enabling the optimizer to take consistent step sizes across the loss landscape. This leads to a more stable search along different dimensions.
We now formalize this optimization process. Let $f_{\mathcal{T}}:\mathcal{V}\rightarrow\mathcal{Z}$ denote the image normalization. Given a clean image $\mathbf{x}\in\mathcal{V}$, where pixel values are bounded within $[0,1]$, and a perturbation constraint mask $\mathbf{r}$, we can first map these values into space $\mathcal{Z}$ for subsequent search:
$\begin{gathered}\mathbf{Z}=f_{\mathcal{T}}(\mathbf{x}),\quad\Delta=f_{\mathcal{T}}(\delta),\quad\mathbf{R}=f_{\mathcal{T}}(\mathbf{r})\\ \underline{\mathbf{Z}}=f_{\mathcal{T}}(\mathbf{0}),\quad\overline{\mathbf{Z}}=f_{\mathcal{T}}(\mathbf{1})\end{gathered}$ (3)
Then, we can focus on the model inference stage $f_{\mathcal{M}}$ and reformulate the optimization objective in Eq.(1) as follows:
$\arg\min_{\Delta}\;\mathbb{E}_{\mathcal{D}}\left\{\mathcal{L}\left(f_{\mathcal{M}}\left(\mathcal{A}\left(\mathbf{Z}^{(c)},\mathbf{R}\right),\mathbf{p}\right),\mathbf{t}^{(c)}\right)\right\}\\$ (4)
with a new constraint function:
$\mathcal{A}\left(\mathbf{Z},\Delta\right)=\left[\mathbf{Z}\cdot\left(\mathbf{1}-\mathbf{R}\right)+\Delta\cdot\mathbf{R}\right]_{\underline{\mathbf{Z}}}^{\overline{\mathbf{Z}}}$ (5)
According to Eq. (4), we can directly search for feasible solutions $\Delta$ within the normalized space. For each semantic clustering $c$, this objective binds the normalized inputs $\mathbf{Z}^{(c)}$ with their corresponding targets $\mathbf{t}^{(c)}$.
After completing the optimization, we obtain the optimal perturbation in the post-normalized space $\Delta^{*}$. We can manually construct an approximately inverse transformation $f_{\mathcal{T}}^{-1}:\mathcal{Z}\rightarrow\mathcal{V}$, with a slight abuse of notation, to recover the adversarial perturbation $\delta$ in the pixel space.
$\delta^{*}=f_{\mathcal{T}}^{-1}(\Delta^{*})$ (6)
Since the perturbation constraints are integrated into the optimization process, the resulting $\delta^{*}$ inherently satisfies the attacker’s requirements.
3.3 Semantic Separation Optimization
Recall that the objective of SAUPs is to mislead MLLMs into generating a predefined target $\mathbf{t}^{(c)}$ corresponding to the class $c$ of a given image. This target $\mathbf{t}^{(c)}$ typically consists of multiple tokens, which can also be denoted as $t^{(c)}_{1:N}$.
[…]
This motivates us to design a proper loss function that not only enhances the alignment between inputs and corresponding targets, but also separates the target labels from each other. To this end, we devise the semantic separation optimization, leveraging cross-entropy loss and margin-based loss. The overall loss function is formulated as follows:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{Margin}}$ (7)
where $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{Margin}}$ denote the cross-entropy loss and the margin loss, respectively. The $\lambda$ is a weighting factor that balances the contributions of the two losses.
Cross-Entropy Guidance for Targets Generation. The cross-entropy loss is a standard approach for inducing specific outputs from MLLMs, inherently enhancing the alignment between inputs and targets within the same semantic cluster. It is defined as follows:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}\log P_{\mathcal{M}}\left(t_{i}^{(c)}\mid t_{<i}^{(c)},\mathcal{A}\left(\mathbf{Z}^{\left(c\right)},\Delta\right),\mathbf{p}\right)$ (8)
where $P_{\mathcal{M}}$ refers to the probability of the token sequence conditioned on the representation of the normalized space.
Inter-Class Discrimination via Margin Loss. To encourage semantic-aware behavior and prevent target obfuscation, we introduce a margin loss that explicitly enforces separation between different targets.
Given an image with the semantic class $c$ and its corresponding target $\mathbf{t}^{(c)}$, we compute the margin loss by measuring the difference in confidence between its correct target $\mathbf{t}^{(c)}$ and all other targets $\mathbf{t}^{(\neg c)}$. These differences are then clipped at zero, and their average is taken as the loss value.
For brevity, we denote $P_{\mathcal{M}}\left(\mathbf{t}\right)$ as a shorthand for the conditional probability $P_{\mathcal{M}}(\mathbf{t}\mid{\mathcal{A}\left(\mathbf{Z},\Delta\right)},\mathbf{p})$. Accordingly, the margin loss can be formally given by:
$\mathcal{L}_{\text{Margin}}=\mathbb{E}_{i\neq c}\left\{\left\lfloor m-P_{\mathcal{M}}\left(\mathbf{t}^{\left(c\right)}\right)+P_{\mathcal{M}}\left(\mathbf{t}^{\left(i\right)}\right)\right\rfloor_{0}\right\}$ (9)
where $m>0$ is a predefined margin enforcing a minimum separation, and $\lfloor\cdot\rfloor_{0}=\max(0,\cdot)$.
4 Latent Space Analysis
@@
In this section, we analyze the latent-space features of images perturbed by SAUPs and reveal three observations. Based on these observations, we propose an explanation for the semantic-aware ability of SAUPs, that is, why a perturbation can discern the semantics of images and assign different targets accordingly.

### v2

# Semantic Router: On the Feasibility of Hijacking MLLMs via a Single Adversarial Perturbation

@@ Abstract
Multimodal Large Language Models (MLLMs) are increasingly deployed in stateless systems, such as autonomous driving and robotics.
This paper investigates a novel threat: Semantic-Aware Hijacking. We explore the feasibility of hijacking multiple stateless decisions simultaneously using a single universal perturbation.
We introduce the Semantic-Aware Universal Perturbation (SAUP), which acts as a semantic router, “actively” perceiving input semantics and routing them to distinct, attacker-defined targets.
To achieve this, we conduct theoretical and empirical analysis on the geometric properties in the latent space. Guided by these insights, we propose the Semantic-Oriented (SORT) optimization strategy and annotate a new dataset with fine-grained semantics to evaluate performance. Extensive experiments on three representative MLLMs demonstrate the fundamental feasibility of this attack, achieving a 66% attack success rate over five targets using a single frame against Qwen.
Image Hijacks, Adversarial Attacks, Multimodal Large Language Models
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) (Liu et al., 2023; Bai et al., 2023; DeepMind, 2025) have emerged as the core perception and reasoning modules for agent applications. In real-world deployments such as autonomous driving (Yang et al., 2024) and robotic manipulation (Kim et al., 2024), the MLLM serves as an “atomic unit” of perception and action. At each time step, the model processes the current frame with a prompt to generate an immediate action, without retaining historical context. For instance, OpenVLA (Kim et al., 2024) creates a new context at each timestep, taking the current frame and an instruction as input, and continuously triggering independent actions. While these atomic decisions are stateless, their sequential accumulation dictates the agent’s physical trajectory.
Figure 1: Illustration of the threat model of semantic-aware hijacking. As the MLLM processes the image sequences, the perturbation acts as a semantic router, forcing the model to output the attacker’s predefined targets based on the input semantics, cumulatively guiding the vehicle to a predefined destination.
This paper focuses on these stateless decisions and investigates a novel and more severe threat: Semantic-Aware Hijacking. We aim to craft a universal adversarial perturbation that acts as a “semantic router,” which forces the model to output distinct, predefined target actions conditioned solely on the visual semantics of the current frame. As illustrated in Figure 1, an adversary can paste an adversarial frame that “actively” senses input semantics (e.g., an intersection vs. a roundabout) and steers the model to generate corresponding targets. We term such perturbations Semantic-Aware Universal Perturbations (SAUPs).
Generating SAUPs imposes a severe geometric constraint on the optimization process. Our theoretical analysis (in Section 3) reveals that mapping semantic inputs to diverse targets requires the perturbation to expand subtle semantic differences into output discrepancies. To overcome optimization challenges, we propose the Semantic-ORienTed (SORT) optimization algorithm. SORT synergizes Normalized Space Optimization to construct a stable optimization manifold, and Semantic Separation Optimization to explicitly maximize the expansion ability.
We annotate the Real Image Sequence Trajectories (RIST) dataset, which serves as a simplified threat model, and enables a quantitative evaluation at varying semantic granularities. Unlike standard classification benchmarks (Deng et al., 2009), RIST targets fine-grained semantics for autonomous driving and robotic scenarios.
@@
We conduct extensive experiments across three MLLMs (Llava-1.5, Qwen2.5-VL, and InternVL3) and two datasets with varying semantic granularities (ImageNet and RIST). Our results validate the fundamental feasibility of the proposed attack. Notably, one perturbation is sufficient to mislead Qwen, achieving success rates of 93%, 77%, 61%, and 66% under 2, 3, 4, and 5 targets, respectively. Code is available at https://anonymous.4open.science/r/semantic-aware-hijacking.
Our main contributions are summarized as follows:
  • We demonstrate the feasibility of hijacking stateless decisions via a single perturbation, and analyze the perturbation in the latent space both theoretically and empirically.
@@ 2.1 Adversarial Perturbations
Adversarial perturbations, as pioneered by (Szegedy et al., 2014; Goodfellow et al., 2014), demonstrate that imperceptible noise added to input data can significantly alter model predictions. A pivotal advancement is the Universal Adversarial Perturbations (UAPs) (Moosavi-Dezfooli et al., 2017; Wang et al., 2023), which are input-agnostic and remain effective across diverse samples. Subsequent works have explored more complex settings: MultiAttack (Fort, 2023) shows that a perturbation can map various inputs to multiple targets, but it lacks universality and remains effective only within the training set. Multi-Target UAP (Custode and Iacca, 2021) seeks to find multiple perturbations simultaneously, yet each one still directs one target.
The primary contribution of this paper is the achievement of a many-to-many UAP. Unlike prior works, this perturbation acts as a “semantic router”, mapping inputs with distinct semantics to different targets. We provide an extended discussion in Appendix F.
@@ 2.2 Vulnerability of MLLMs
The rapid evolution of MLLMs (Liu et al., 2023; Bai et al., 2023) has introduced significant security concerns. Recent studies (Qi et al., 2024; Shayegani et al., 2024; Zhang et al., 2024; Ying et al., 2024) reveal that MLLMs inherit vulnerabilities to adversarial attacks. Existing studies can be generally categorized into two types:
Semantic misleading (Zhao et al., 2023; Liu et al., 2024; Xu et al., 2025) focuses on causing the model to misinterpret the visual content. These attacks typically target the vision encoder, feeding corrupted representations to the downstream language model to induce incorrect reasoning outcomes. For instance, C-PGC (Fang et al., 2024) proposes a UAP against vision-language models that requires manipulation of both image and text modalities.
Image hijacking (Luo et al., 2024; Lu et al., 2024; Bailey et al., 2023) seeks token-level control over the model’s output. These attacks often utilize end-to-end optimization. Such methods can force MLLMs to execute malicious payloads, such as generating bash commands or bypassing safety filters in downstream tasks (Bailey et al., 2023).
3 Geometric Mechanics of SAUPs
3.1 Problem Formulation and Intuition
Threat Model. This paper aims to demonstrate the fundamental feasibility of hijacking MLLMs’ stateless decisions via a single and universal perturbation. Consequently, our analysis is conducted under a digital white-box setting to verify the latent space behavior, prioritizing the understanding of the attack mechanism over physical deployment.
We focus on stateless decisions, where the model functions as an atomic unit without retaining historical memory (e.g., OpenVLA (Kim et al., 2024)). In this paradigm, hijacking the atomic decision at each step effectively dictates the agent’s cumulative physical trajectory.
Definition 3.1.
Stateless decisions is a sequence of independent actions $\{y_{1},\dots,y_{K}\}$ indexed by time $k\in[1,K]$. At each step $k$, the MLLM, denoted as $\mathcal{M}$, generates an action $y_{k}$ based solely on the current visual observation $x_{k}\in\mathbb{R}^{d}$ and the textual prompt $p_{k}$, formulated as:
$y_{k}=\mathcal{M}(x_{k},p_{k})$ (1)
Note that the generation of $y_{k}$ is independent of previous states $\{(x_{i},p_{i},y_{i})\}_{i<k}$, meaning the model retains no historical context between steps.
The adversary seeks a single perturbation $\delta$ to hijack multiple stateless decisions simultaneously. Specifically, when the perturbation is added to the visual input $x^{(c)}$ from a specific semantic category $c$, it compels the model to output a predefined target label $t^{(c)}$. We refer to such perturbations as Semantic-Aware Universal Perturbation (SAUP).
$\mathcal{M}(x^{(c)}+\delta,p^{(c)})\to t^{(c)},\quad\forall c\in\{1,\dots,C\}$ (2)
where $C$ denotes the number of distinct semantic classes.
Figure 2: Illustration of the geometric mechanism of the Semantic-Aware phenomenon. Given three image sets $\{x^{(c)}\}$ with distinct semantic contents, and an all-zero pixel image serving as an anchor. The SAUP maps all image features to a distant region in the latent space, while the respective semantics cause slight deflections, guiding them towards alignment with predefined targets.
Intuitive Mechanism.
Before diving into the geometric derivation, we provide an intuitive explanation of how a single perturbation $\delta$ achieves this complex mapping, as illustrated in Figure 2. The mechanism relies on two key geometric effects in the latent space:
Dominant Shift: When the perturbation is introduced to images, its features become dominant, shifting input clusters away from their original manifolds toward the adversarial subspace. In this subspace, decision boundaries are highly dense and congested (Liu et al., 2016).
Semantic Deflection: Once shifted, the perturbation leverages the intrinsic semantic differences between inputs to deflect features towards distinct directions, eventually aligning them with predefined targets.
3.2 Geometric Decomposition in Latent Space
To understand how visual inputs define textual outputs, we decompose the MLLM $\mathcal{M}$ into a vision encoder $\phi:\mathcal{X}\to\mathcal{Z}$ mapping images to embeddings of latent space, and a backbone decoder model $\mathcal{D}:\mathcal{Z},\mathcal{P}\to\mathcal{Y}$ .
A key challenge in formulation is that the target $t^{(c)}$ is a text sequence, while the perturbation operates in the visual input space. To facilitate a theoretical analysis, we introduce the concept of Proxy Target Embeddings.
Assumption 3.2.
For a given target response $t^{(c)}$ and prompt $p$, we assume there exists a corresponding visual embedding $\hat{z}^{(c)}\in\mathcal{Z}$, which can induce the decoder $D$ generate $t^{(c)}$ (i.e., $D(\hat{z}^{(c)},p)\to t^{(c)}$). We term this visual embedding $\hat{z}^{(c)}$ as the Proxy Target Embedding [^1] While we denote $\phi$ as the vision encoder for simplicity, the concept of Proxy Target Embedding can be generalized to any intermediate layer where semantic alignment occurs..
Under this assumption, the goal of the perturbation $\delta$ is to map the input set $\{x^{(c)}\}$ to the corresponding set of proxy target embeddings $\{\hat{z}^{(c)}\}$. We analyze this mapping mechanism via a first-order Taylor expansion of the vision encoder $\phi$ around the perturbation $\delta$.
For an input image $x^{(c)}$ belonging to class $c$, the feature representation of the perturbed image $z^{(c)}=\phi(x^{(c)}+\delta)$ can be approximated as:
$z^{(c)}\approx\underbrace{\phi(\delta)}_{\text{Dominant Shift }\mu_{\delta}}+\underbrace{J_{\delta}\cdot x^{(c)}}_{\text{Semantic Deflection }\nu_{c}}$ (3)
where $J_{\delta}=\frac{\partial\phi(x)}{\partial x}\big|_{x=\delta}$ is the Jacobian matrix evaluated at the perturbation point, $\phi(\delta)$ represents a Dominant Shift, and $\nu_{c}=J_{\delta}x^{(c)}$ represents the Semantic Deflection.
Remark 3.3.
The optimization of SAUP is geometrically equivalent to searching for a perturbation $\delta$ such that the perturbed embedding $z^{(c)}$ aligns with the corresponding proxy target embedding $\hat{z}^{(c)}$. Specifically, the local linear projection $J_{\delta}$ serves to extract the intrinsic semantics of $x^{(c)}$ and map them toward class-specific targets.
3.3 Semantic Separability Potency and Bounds
Recall from Equation 3 that the perturbation $\delta$ should preserve and amplify the unique semantic of $x^{(c)}$. We define the Deflection Strength as the magnitude of its deviation from the anchor point $\phi(\delta)$:
$\mathcal{S}(\delta,x^{(c)})=\|\phi(x^{(c)}+\delta)-\phi(\delta)\|_{2}$ (4)
Using the first-order Taylor approximation [^2] The derived bound holds approximately by neglecting the higher-order terms of the Taylor series. While it is not strictly exact, the qualitative conclusion still holds., this strength is determined by the local geometry at $\delta$:
$\mathcal{S}(\delta,x^{(c)})\approx\|J_{\delta}x^{(c)}\|_{2}\leq\|J_{\delta}\|_{2}\cdot\|x^{(c)}\|_{2}$ (5)
where $\|J_{\delta}\|_{2}$ is the spectral norm of the Jacobian matrix at the perturbation point. Here, $\|J_{\delta}\|_{2}$ serves as the amplification factor, quantifying the capability of semantic deflection.
To hijack multiple decisions, the adversary requires more than just sensitivity; they require separability. For any pair of distinct classes $i$ and $j$, the perturbation must identify the semantic difference between inputs $x^{(i)}$ and $x^{(j)}$, and align with corresponding proxy target embeddings $\hat{x}^{(i)}$ and $\hat{x}^{(j)}$.
$\|\phi(x^{(i)}+\delta)-\phi(x^{(j)}+\delta)\|_{2}\approx\|\hat{z}^{(i)}-\hat{z}^{(j)}\|_{2}$ (6)
Substituting the linearization2, the achievable separation in the latent space is governed by:
$\|\hat{z}^{(i)}-\hat{z}^{(j)}\|_{2}\leq\underbrace{\|J_{\delta}\|_{2}}_{\text{Potency}}\cdot\underbrace{\|x^{(i)}-x^{(j)}\|_{2}}_{\text{Input Distance}}$ (7)
Let $L=\sup_{z}\|J_{z}\|_{2}$ be the global Lipschitz constant of the vision encoder. We define the Required Expansion Ratio $\rho_{ij}$ for a input-target pair as:
$\rho_{ij}=\frac{\|\hat{z}^{(i)}-\hat{z}^{(j)}\|_{2}}{\|x^{(i)}-x^{(j)}\|_{2}}$ (8)
Theorem 3.4.
Let $\rho_{\max}=\max_{i\neq j}\rho_{ij}$ be the Maximum Required Expansion Ratio. If $\rho_{\max}>L$, then no perturbation $\delta$ exists that can successfully map the input set $\{x^{(c)}\}$ to the target set $\{\hat{z}^{(c)}\}$.
Remark 3.5.
This theorem highlights the difficulty of fine-grained semantics. As the images becomes similar (e.g., images captured by the same dashboard camera), the input distance $\|x^{(i)}-x^{(j)}\|_{2}\to 0$, causing $\rho_{\max}\to\infty$, which makes it increasingly likely to violate the bound $L$.
4 Methodology: SORT
Guided by the theoretical analysis in Section 3, we propose the Semantic-ORienTed (SORT) algorithm. As illustrated in Figure 3, SORT incorporates a normalized space optimization to stabilize gradients, and a semantic separation strategy to explicitly maximize the deflection potency $J_{\delta}$.
@@ 4 Methodology: SORT
4.1 Normalized Space Optimization (NSO)
To facilitate the addition of perturbations to images, existing studies (Bailey et al., 2023; Lu et al., 2024; Wang and He, 2021; Vo et al., 2022) conduct their optimization directly in the pixel space ([0, 1]). This hinders the precise adjustment of the Jacobian $J_{\delta}$. To ensure stable updates, we perform a change of variables into the normalized space.
Let $\Psi:\mathcal{X}\to\mathbb{R}^{d}$ be the normalization function (e.g., standardization). We define a trainable variable $\Delta$ in the normalized space. The perturbation $\delta$ in pixel space is recovered via the inverse transformation:
$\delta=\Psi^{-1}(\Delta)$ (9)
By optimizing $\Delta$ directly, we precondition the optimization landscape, ensuring consistent step sizes and avoiding the plateauing loss observed in pixel-space constraints.
4.2 Semantic Separation Optimization (SSO)
To search for a feasible perturbation, we design a hybrid objective that combines cross-entropy loss $\mathcal{L}_{CE}$ and a margin-based loss $\mathcal{L}_{Margin}$. The cross-entropy loss ensures basic alignment with the target, while the margin-based loss explicitly forces input embeddings apart.
Recall Equation 6 that the optimization must effectively expand the feature distance $\|\phi(x^{(i)}+\delta)-\phi(x^{(j)}+\delta)\|_{2}$ to align with the distance between proxy target embeddings $\|\hat{z}^{(i)}-\hat{z}^{(j)}\|_{2}$. Therefore, we employ the margin loss to explicitly drive the perturbed features apart by maximizing the separation in the output space.
Let $P(\cdot|x,p)$ denote the output confidence of the MLLM, conditioned on the visual input $x$ and textual prompt $p$. We have the Confidence Gap $\Delta P_{cj}$ between the corresponding target $t^{(c)}$ and a competing target $t^{(j)}$:
$\Delta P_{cj}=P(t^{(c)}|x^{(c)}+\delta,p)-P(t^{(j)}|x^{(c)}+\delta,p)$ (10)
To enforce the separation, we maximize this gap until it exceeds a specified margin $m$:
$\mathcal{L}_{Margin}=\mathbb{E}_{j\neq c}\left[\max\left(0,m-\Delta P_{cj}\right)\right]$ (11)
Geometrically, Minimizing $\mathcal{L}_{Margin}$ requires the perturbed feature $\phi(x^{(c)}+\delta)$ to move closer to $\hat{z}^{(c)}$ than to $\hat{z}^{(j)}$. The final objective is:
$\mathcal{L}_{Total}=\mathcal{L}_{CE}+\lambda\cdot\mathcal{L}_{Margin}$ (12)
where $\lambda$ is a balancing hyperparameter that controls the trade-off between target alignment and semantic separation.
5 Validation of Geometric Mechanics
In this section, we empirically validate the theoretical framework proposed in Section 3 by analyzing the latent space representations of the perturbed images. Our goal is to verify the decomposition of the perturbed feature $z^{(c)}$ into the Dominant Shift ($\mu_{\delta}$) and the Semantic Deflection ($\nu_{c}$), as formulated in Equation 3.